Scooby Doo is initially at rest when he sees a box of Scooby Snacks 55 m up ahead. If Scooby Doo accelerates at 0.44 m/s2, how long does it take him to reach the box of Scooby Snacks? 


1
Expert's answer
2021-01-13T11:37:52-0500

The distance travelled under a constant acceleration from rest is:


"d = \dfrac{at^2}{2}"

where "d = 55m", "a = 0.44m\/s^2", and "t" is the time of motion. Expressing the time, obtain:


"t = \sqrt{\dfrac{2d}{a}}\\
t = \sqrt{\dfrac{2\cdot 55}{0.44}} \approx 15.81s"

Answer. 15.81s.